### v2

@@ 3 Numerical model
Once the equilibrium is reached a snapshot of the sample is taken. This allows us to track the evolution of different trackers presented in Section 4. Those steps are repeated until all the bonds are broken. Notice that a bond can break because of the dissolution (the surface is null or negative) or because of the loading (criteria presented Equation 22 are reached).
@@
Bond dissolution $0.02\,e^{m_{log}}$ $0.05\,e^{m_{log}}$ (when $95$ of the initial number of bonds are broken)
@@
Looking at Table 2, the Young modulus $E_{m}$ appears to depend on the degree of cementation. An update of the Young modulus (and so, of the different contact springs $K_{n}$, $K_{s}$, $K_{r}$, $K_{t}$) is done during the simulation following Equation 24. The influence of this stiffness reduction will be also investigated in this paper in Section 4.2.
@@
where $E_{m}$ is the current Young modulus used in the simulation, $E_{m}^{cementation}$ is the initial Young modulus depending on the initial degree of cementation, $E_{m}^{untreated}$ is the Young modulus without cementation (here $80$ MPa), and $\xi$ is the debonding factor defined as the ratio of the bonds dissolved (by dissolution or by loading) over the initial number of the bonds.
The time step $dt$ must verify the P-wave critical time step condition [35] defined as:
$dt<dt_{crit}=\frac{\text{min}(R)}{\sqrt{E_{m}/\rho}}$ (25)
where $R$ is the radius of the grains, $E_{m}$ is the Young modulus and $\rho$ is the density of the material. Here, a safety factor is considered, $dt=0.6\,dt_{crit}$. As the Young modulus $E_{m}$ evolves with the dissolution of the bonds, an update of this time step $dt$ is also done.
@@ 4 Results
The results of all the simulations performed (plus some complementary investigations) are presented in the following subsections. Two main trackers are used:
  • 1.
@@
Thanks to the DEM, much more data are available to understand the mechanisms operating at the microstructural level as the coordination number, the mode of rupture of the bonds or the mean force transmitted in the contact, among others. They will be used if necessary to understand our results.
@@ 4.1 The existence of an attractor value $k_{0}^{attr}$
The evolutions of $k_{0}$ are gathered in Figure 4 for all the cementations, the three values of $k_{0}^{init}$ and the different confining pressure $P_{confinement}$.
@@
It appears the $k_{0}$ aims to reach an attractor value $k_{0}^{attr}$ with the debonding phenomena. Applying the principle of superposition, it has been highlighted that $k_{0}$ approaches a limit given by $k_{0}^{attr}=\nu/(1-\nu)$ during weathering and stiffness reduction [19]. Considering a window of $[0.2-0.3]$ for the value of $\nu$, the attractor value window is $[0.25-0.43]$.
To understand better this phenomenon, the fabric description of the sample and its evolution are investigated. Three simulations are considered: a 13MB sample with a confining pressure $P_{confinement}=1$ MPa and an initial value of $k_{0}=0.2$, $0.4$ or $0.7$. Even if a contact-based fabric tensor is often used in the literature [37, 38, 39], a normal forces-based fabric tensor $F^{n}_{ij}$ [39, 40, 41] is used here. Indeed, the behavior of the grain in the oedometric condition is dominated by the normal forces and the focus is on mechanical anisotropy. The normal forces-based tensor is defined by the average of the outer product of the normal forces at contacts, see Equation 26.
@@
where $N_{c}$ is the total number of contacts, $c$ is a contact in the sample and $n_{i}$ is the unit vector of the contact direction. A distribution function $f^{n}(n_{i})$ can be defined in Equation 29 [39] and illustrated in Figure 5 at initial conditions and at final conditions for $k_{0}^{init}=0.2$, $0.4$ and $0.7$. Even if the initial value of the $k_{0}$ influences the probability distribution function at the initial step, it appears that the fabric aims to reach an attractor configuration thanks to the grain reorganization. This attractor configuration is computed considering the average of the three final configurations.
@@
where $f_{0}=tr(F^{n}_{ij})$ and $a^{n}_{ij}=15/2\left(F^{n}_{ij}/f_{0}-\delta_{ij}\right)$. This function represents the average normal force on a specific direction [40].
Figure 5: Distribution function $f^{n}(n_{i})$ at initial and attractor conditions. This function represents the average normal force on a specific direction.
It appears the $k_{0}^{init}=0.2$ function (blue) is inside the $k_{0}^{init}=0.4$ function (red) which is inside the $k_{0}^{init}=0.7$ function (green). Indeed, it appears in Figure 9, used later in this work, that the mean force transmitted in contact is larger for the case $k_{0}^{init}=0.7$, then for $k_{0}^{init}=0.4$ and finally for $k_{0}^{init}=0.2$. It is in agreement with the definition of the function $f^{n}$. Moreover, it is highlighted that the larger $k_{0}^{init}$ is, the more similar to a circle $f^{n}$ is, for planes yz and xz. The sphericity means that the sample is more isotropic, in agreement with the values of $k_{0}^{init}$ closer to $1$.
One can notice, on plane xy, the fact that an anisotropy exists. Indeed, because of the initial configuration algorithm, see Figure 2, especially the fact that only one wall is moving to apply $k_{0}^{init}$, one loading direction is favored.
Then, this probability distribution function can be normalized to compare the different organizations. A normalized function $\overline{f^{n}}$ is defined in Equation 30 and shown in Figure 6.
@@
Figure 6: Distribution function $\overline{f^{n}(n_{i})}$ at initial and attractor conditions. This function represents the probability of a contact to exist in a direction weighted by the average normal force in this direction.
The normalization helps to compare the cases. Looking at planes yz and xz, it is clear that the $k_{0}^{init}=0.2$ case must create isotropy ($k_{0}$ will increase), whereas the $k_{0}^{init}=0.4$ or $=0.7$ cases must create anisotropy ($k_{0}$ will decrease). The same observations than sooner can be done on the plane xy, one direction is favorized because of the initial condition algorithm, see Figure 2.
@@
It appears that in planes yz and xz, the maximum values increase and the minimum values decrease for cases $k_{0}^{init}=0.4$ and $=0.7$. This distancing of the extrema reveals the creation of anisotropy. On the opposite, the maximum values decrease and the minimum values increase for the case $k_{0}^{init}=0.2$, revealing the creation of isotropy (closing of the extrema). Those observations are in agreement with the evolution of the $k_{0}$ (proxy of the isotropy/anisotropy of the sample). In plane xy, the extrema are closing for the three cases (creation of isotropy). Indeed, the setup presented here does not induce anisotropy on this specific plane. This initial anisotropy is generated by the initial configuration algorithm, see Figure 2.
4.2 Influence of the Young modulus reduction assumption
In this section, the Young modulus reduction assumption during dissolution, described in Equation 24, is discussed and the $k_{0}$ evolution mechanisms are investigated. The same simulations are run without the Young modulus reduction assumption (the Young modulus stays constant during the debonding of the sample). The results are presented in Figure 8.
@@ 4.2 Influence of the Young modulus reduction assumption
In the case $k_{0}^{init}\leq k_{0}^{attr}$ (here an example with $k_{0}^{init}=0.2$), illustrated in Figure 8a, it appears that a grain reorganization occurs ($k_{0}$ evolves during the dissolution) even if the Young modulus stays constant. It appears in Figure 8b that there is no grain reorganization ($k_{0}$ stays constant during the dissolution) if the Young modulus stays constant and if $k_{0}^{init}\geq k_{0}^{attr}$ (here an example with $k_{0}^{init}=0.7$). In this kind of configuration, the reorganization of the grains is due to the softening of the material (Young modulus reduction). Two mechanisms seem to appear, depending on the state of stress.
This phenomenon is also highlighted by considering the evolution of the mean force and the coordination number (proxy of the grain organization). Three simulations are considered: a 13MB sample with a confining pressure $P_{confinement}=1$ MPa and an initial value of $k_{0}=0.2$, $0.4$ or $0.7$. The results are shown in Figure 9.
@@
The existence of the two mechanisms is emphasized by the fact that the $k_{0}^{init}=0.2$ mean force evolution is different than the $k_{0}^{init}=0.4$ or $=0.7$ mean force evolutions. The first one stays at a constant value ($\approx 0.02$ N) until a threshold value ($\xi\approx 0.6$) for the debonding is reached. The latter decreases linearly from the beginning of the simulation. A second evolution (sharper and less linear) can be spotted after a threshold debonding ($\xi\approx 0.8$). In the same way, it appears the $k_{0}^{init}=0.2$ coordination number aims to reduce at the start of the simulation whereas the $k_{0}^{init}=0.4$ and $=0.7$ coordination numbers increase. This observation reinforced the idea that two mechanisms exist, depending on the state of stress.
@@
4.3 Influence of the initial $k_{0}$ on the grain reorganization mechanism
As emphasized in Section 4.2, the grain reorganization mechanism (evolution of the $k_{0}$) seems to depend on the state of stress. The distribution of the mechanisms of bond rupture presented in Figure 10 emphasizes this phenomenon. The bonds can break in two modes: i) by dissolution (if the bond surface reaches a null or a negative value) or ii) by loading (if criteria presented in Equation 22 are reached). It appears that for the same cementation and the same confining pressure $P_{confinement}$, the percentage of rupture by loading is larger for $k_{0}^{init}\leq k_{0}^{attr}$ than for $k_{0}^{init}\geq k_{0}^{attr}$. To understand this difference, the chain forces are sorted into two kinds: unstable and stable organization. An unstable chain force stays in place thanks to the cementation at the contacts. Once the cement is partially dissolved, the chain force collapses, breaking the reduced but existing cemented contacts. On the opposite, a stable chain force stays in place thanks to the particle-particle contact. If the cemented contacts are dissolved, the chain force does not collapse. Considering $k_{0}^{init}\leq k_{0}^{attr}$, it has been established that grain reorganization occurs only if enough cemented contacts are reduced. Indeed, the unstable chain forces are predominant in this case, see the $k_{0}$ evolution when the Young modulus $E_{m}$ stays constant during the debonding, Figure 8a. Once the cemented contacts are reduced, the cementation at contacts breaks by mechanical loading. Considering $k_{0}^{init}\geq k_{0}^{attr}$, it has been established that grain reorganization occurs directly (no cemented contact reduction is needed). Indeed, the stable chain forces are predominant in this case, see the constant $k_{0}$ when the Young modulus $E_{m}$ stays constant during the debonding, Figure 8b. The cemented contacts break with a chemical rupture as they are not sollicitated.

### v3

@@ 3 Numerical model
It is crucial to provide a commentary on the concept of grain reorganization. In the configuration under investigation, the positions of the particles remain largely unchanged. Indeed, the sample is dense and in oedometric conditions. Furthermore, it is assumed that no grain breakage occurs. Nevertheless, some fluctuations in the organization do occur, resulting in an evolution in the chain force structure.
Once the equilibrium is reached, a snapshot of the sample is taken. This enables the tracking of the evolution of the various trackers presented in Section 4. These steps are repeated until all the bonds have disappeared. It should be noted that a bond can disappear because of the dissolution (the surface is null or negative) or the mechanical loading (criteria presented in Equation 23 are reached).
@@
Bond dissolution $0.02\,e^{m_{log}}$ ($0.05\,e^{m_{log}}$ when $95$ of the initial number of bonds are broken)
@@
To explore the effect of the cementation, five distinct initial cementations, exhibiting disparate microscopic parameters, are considered. During the debonding process, the bond surfaces $A_{b}$ are reduced, and the sample aims to reach the untreated granular state. The reduction of the bond surfaces results in a reduction in the strength of the cemented contacts, see Equation 23. Such weakening may ultimately result in the mechanical rupture of the bond. Moreover, in the event that the bond remains intact and its surface area reaches $A_{b}=0m^{2}$, it may also undergo a chemical rupture. These ruptures weaken the sample, and the current state during the debonding phenomenon is interpolated from the only known states: the initial cemented state and the final unbonded state, see Figure 6.
Figure 6: During the debonding phenomenon, the current state is interpolated from the initial (cemented) and final (unbonded) states.
Table 2 highlights the fact that the Young modulus $E_{m}$ exhibits variation between the unbonded and bonded states. To ensure the sample is representative, and because of the data available in [53], an interpolation between the two known states is performed to characterize the current state. This approach is commonly applied in the literature to model chemical damage at the macroscale [35, 68, 69, 70]. In particular, the Young modulus (and thus, the different contact springs $K_{n}$, $K_{s}$, $K_{r}$, $K_{t}$) is updated following Equation 25. This assumption results in the dependency of the contact scale on the sample scale through the debonding factor $\xi$. In this investigation, a linear approximation is employed as it represents the most straightforward relation available for a system comprising only two known states. Should supplementary data become available, this relation can be readily modified in accordance with the new information. The impact of this stiffness reduction will be examined in this paper in Section 4.2 and a discussion will be presented in Section 5.2.
@@
where $E_{m}$ is the current Young modulus used in the simulation, $E_{m}^{cementation}$ is the initial Young modulus depending on the initial degree of cementation, $E_{m}^{untreated}$ is the Young modulus without cementation (here $80$ MPa), and $\xi$ is the debonding factor defined as the ratio of the bonds dissolved (by dissolution or by loading) over the initial number of the bonds. As the time step employed for the temporal integration depends on the Young modulus, see Equation 3, it becomes important to update the time increment $dt$ with the dissolution of the bonds. The relationship between the contact scale ($E_{m}$) and the sample scale ($\xi$) could be a matter of contention. An alternative approach would have been to consider a local surface stiffness $E_{m}(A_{b})$ at each contact, with the stiffness dependent on the bond surface [71, 72]. This approach is explored in Section 5.3. However, this relationship is not available in the data set used [53], and can only be extrapolated. Consequently, the model choice has been to interpolate the current state, see Figure 6. Nevertheless, contact strength depends on the local bond surface area $A_{b}$, see Equation 23.
@@ 4 Results
The results of all the simulations performed (plus some complementary investigations) are presented in the following Subsections. Two main variables are investigated:
@@
Thanks to the DEM, much more data are available to understand the mechanisms operating at the microstructural level such as the coordination number, the mode of rupture of the bonds or the mean force transmitted in the contact, among others. These data will be used if necessary to understand our results.
@@ 4.1 The existence of an attractor value $k_{0}^{attr}$
The evolutions of $k_{0}$ are gathered in Figure 7 for the five initial degrees of cementation, the three values of $k_{0}^{init}$ and the three distinct confining pressures $P_{confinement}$.
@@
It appears the $k_{0}$ aims to reach an attractor value $k_{0}^{attr}$ with the dissolution of the bonds. Applying the principle of superposition, it has been shown analytically by [28] that $k_{0}$ approaches a limit given by $k_{0}^{attr}=\nu/(1-\nu)$ during weathering and stiffness reduction. Considering a window of $[0.2-0.3]$ for the value of $\nu$ [74, 75, 76], the attractor value window is $[0.25-0.43]$, in agreement with the results obtained. This evolution of the $k_{0}$ illustrates the granular reorganization [44, 73]. The sample being dense and under oedometric conditions, the positions of the particles remain mostly unchanged. Nevertheless, these small fluctuations affect widely the distribution of forces between particles and thus the stress state of the sample.
To understand better the attractor configuration, the fabric description of the sample and its evolution are investigated. Three simulations are considered: a 13MB sample with a confining pressure $P_{confinement}=1$ MPa and initial values of $k_{0}=0.2$, $0.4$ or $0.7$. While contact-based fabric tensors are frequently referenced in the literature [77, 78, 79], this study employs a normal forces-based fabric tensor $F^{n}_{ij}$ [79, 80, 81]. Indeed, the normal forces-based fabric tensor underlines information on the fabric anisotropy, weighted by the norm of the normal force transmitted at the contacts. Being dense and under oedometric conditions, the normal forces are dominant in the stress transmission. The debonding process results in an evolution of chain forces, even if the position of the grains remains similar. The primary cause of this evolution can be attributed to the modification of normal forces. To quantify this variation, a normal forces-based tensor is defined as the average of the outer product of the normal forces at contacts, see Equation 26.
@@
where $N_{c}$ is the total number of contacts, $c$ is a contact in the sample and $n_{i}$ is the unit vector of the contact direction. A distribution function $f^{n}(n_{i})$ can be defined in Equation 29 [79] and illustrated in Figure 8 at initial conditions for $k_{0}^{init}=0.2$ (blue), $0.4$ (red) and $0.7$ (green) and at the attractor conditions (dotted). Even if the initial value of the $k_{0}$ influences the probability distribution function at the initial state, it appears that the fabric aims to reach an attractor configuration thanks to the granular reorganization. This attractor configuration is computed considering the average of the three final configurations (individually similar to the attractor).
@@
where $f_{0}=tr(F^{n}_{ij})$ and $a^{n}_{ij}=15/2\left(F^{n}_{ij}/f_{0}-\delta_{ij}\right)$. This function represents the average normal force in a specific direction [80].
Figure 8: Distribution function $f^{n}(n_{i})$ at initial and attractor conditions (dotted). This function represents the average normal force on a specific direction.
It can be observed that the function $k_{0}^{init}=0.2$ (depicted in blue) is situated within the function $k_{0}^{init}=0.4$ (depicted in red), which in turn is situated within the function $k_{0}^{init}=0.7$ (depicted in green). As illustrated in Figure 12, the mean force transmitted in contact is found to be larger for the case $k_{0}^{init}=0.7$ than for either $k_{0}^{init}=0.4$ or $k_{0}^{init}=0.2$. This is consistent with the definition of the function $f^{n}$, representing the average normal force in a specific direction. Furthermore, it is observed that as $k_{0}^{init}$ increases, the distribution of $f^{n}$ for the planes yz and xz approaches a circle. The sphericity of $f^{n}$ indicates that the distribution of the forces is more isotropic, which is consistent with the values of $k_{0}^{init}$ closer to $1$. This tendency is even clearer in Figure 9, where the distribution function is normalized.
One can notice the fact that an anisotropy exists on plane xy. Indeed, the initial configuration algorithm consists of moving only one wall to apply $k_{0}^{init}$, favoring one loading direction, see Figure 4. However, this anisotropy remains insignificant.
Then, this probability distribution function is normalized to compare the different organizations. A normalized function $\overline{f^{n}}$ is defined in Equation 30 and shown in Figure 9.
@@
Figure 9: Distribution function $\overline{f^{n}}(n_{i})$ at initial and attractor conditions (dotted). This function represents the probability of a contact to exist in a direction weighted by the average normal force in this direction.
The normalization enables us to compare the cases. Looking at planes yz and xz, it is clear that the $k_{0}^{init}=0.2$ (blue) case must create isotropy ($k_{0}$ will increase), whereas the $k_{0}^{init}=0.4$ (red) or $=0.7$ (green) cases must create anisotropy ($k_{0}$ will decrease). As a reminder, an isotropic state of stress is defined as $k_{0}=1$, and it is depicted as a perfect circle. The same observations can be done on the plane xy, one direction is favored because of the initial condition algorithm, see Figure 4. However, this anisotropy remains insignificant.
@@
It appears that in planes yz and xz, the maximum values increase and the minimum values decrease for cases $k_{0}^{init}=0.4$ and $=0.7$. This distancing of the extrema reveals the creation of anisotropy. On the opposite, the maximum values decrease and the minimum values increase for the case $k_{0}^{init}=0.2$, revealing the creation of isotropy (closing of the extrema). These observations are in agreement with the evolution of the $k_{0}$, proxy of the isotropy/anisotropy of the sample [44, 73]. In plane xy, the extrema are closing for the three cases (creation of isotropy). The initial anisotropy is generated by the initial configuration algorithm, see Figure 4, and it is diminished with the debonding phenomenon.
It is crucial to acknowledge that $k_{0}$ is computed in this study from the stress applied on a single lateral wall. One might inquire as to how the stress states evolve in the minor and intermediate principal directions. Figures 8, 9 and 10 show that anisotropy exists initially on the plane xy (especially for $k_{0}^{init}=0.4$ and $0.7$). This is because of the initialization process, see Figure 4. Indeed, only a single wall (with the normal in the x-direction) is moving to control $k_{0}$. Minor and intermediate stresses are different, even if they stay similar. Minor stress is in the direction used to compute $k_{0}$. With the debonding the anisotropy on this plane disappears, see the attractor line (dotted), and minor and intermediate stresses aim to reach the same value.
4.2 The existence of two $k_{0}$ evolution mechanisms
This Section addresses the issue of the reduction in Young modulus during the dissolution process, see Equation 25. It also investigates the mechanisms of the $k_{0}$ evolution. The identical simulations conducted in the preceding Section are now executed without the assumption of a reduction in the Young modulus (the Young modulus is held constant throughout the debonding process). The results are presented in Figure 11.
@@ 4.2 The existence of two $k_{0}$ evolution mechanisms
In the case $k_{0}^{init}\leq k_{0}^{attr}$, the data illustrated in Figure 11a appears to demonstrate that a grain reorganization occurs ($k_{0}$ evolves) even when the Young modulus remains constant. On the other hand, Figure 11b illustrates that grain reorganization does not occur ($k_{0}$ remains constant) when the Young modulus remains constant for $k_{0}^{init}\geq k_{0}^{attr}$. In this configuration, the reorganization of the grains is a consequence of the reduction in the material’s Young modulus. Two distinct mechanisms appear to emerge, contingent on the state of stress.
This phenomenon is also highlighted by considering the evolution of the mean force and the coordination number (proxy of the grain organization). Three simulations are considered: a 13MB sample with a confining pressure $P_{confinement}=1$ MPa and initial values of $k_{0}=0.2$, $0.4$ or $0.7$. The results are shown in Figure 12.
@@
The distinction between the two mechanisms is underscored by the observation that the mean force evolution corresponding to $k_{0}^{init}=0.2$ exhibits a distinct trajectory compared to the mean force evolutions associated with $k_{0}^{init}=0.4$ or $=0.7$. The first one remains constant ($\approx 0.02$ N), until a threshold value for the debonding is reached ($\xi\approx 0.6$). In contrast, the mean forces associated with $k_{0}^{init}=0.4$ or $=0.7$ decrease in a linear fashion from the outset of the simulation. A second evolution, exhibiting a more pronounced and less linear trend, can be observed following the attainment of a debonding threshold ($\xi\approx 0.8$). Similarly, it seems that the coordination number corresponding to $k_{0}^{init}=0.2$ reduces at the outset of the simulation. Whereas, the coordination numbers corresponding to $k_{0}^{init}=0.4$ and $=0.7$ are intended to increase. This observation served to reinforce the hypothesis that two distinct mechanisms are at play, contingent on the state of stress.
@@
4.3 Influence of the initial stress state on the main grain reorganization mechanism
As emphasized in Section 4.2, the principal grain reorganization mechanism seems to depend on the state of the stress. This reorganization induces an evolution of $k_{0}$, related to the evolution of the chain force [44, 73]. It should be noted that the positions of the particles remain largely unaltered as the sample is dense in oedometric conditions. Nevertheless, the minor fluctuations impact the state of the stress within the sample.
The distribution of the mechanisms of bond rupture, illustrated in Figure 14, highlights the diverse stress reorganization mechanisms. This distribution considers only the bond rupture after the sample initialization described in Section 3. Indeed, as explained in Section 3, this preparation step induces bond breakage before the commencement of the bond dissolution, in a manner analogous to that observed in experiments. The bonds may break in two modes: i) by dissolution (if the bond surface $A_{b}$ reaches a null or a negative value) or ii) by mechanical loading (if criteria presented in Equation 23 are reached). It can be observed that, for a given cementation and confining pressure, the percentage of rupture by loading is greater for $k_{0}^{init}\leq k_{0}^{attr}$ than for $k_{0}^{init}\geq k_{0}^{attr}$. To gain insight into this discrepancy, the chain forces are classified into two categories: the unstable chain force and the stable chain force, see Figure 13.
Figure 13: The granular media comprises unstable and stable chain forces.
An unstable chain force is maintained in position by the cementation at the contacts. Once the cement has undergone partial dissolution, the chain force collapses. It should be noted that while the particle positions remain largely unchanged, minor fluctuations induce the evolution of the chain force, resulting in the breaking of the reduced but existing cemented contacts. These unstable chain forces induce the bond rupture by mechanical loading. In contrast, a stable chain force remains in place due to the particle-particle contact. In the event of the total dissolution of the cemented contacts, the chain force does not collapse. These stable chain forces induce the bond rupture by chemical loading. It has been demonstrated in the previous Section that grain reorganization will only occur in the case $k_{0}^{init}\leq k^{attr}_{0}$ if the cemented contacts are reduced enough. Indeed, the unstable chain forces are predominant in this case. For an illustration of this, one may refer to the $k_{0}$ evolution when the Young modulus $E_{m}$ remains constant during the debonding process, Figure 11a. Even if the Young modulus remains constant during the debonding, $k_{0}$ evolves/granular reorganization occurs. Once the cemented contacts have been reduced, the cementation at the contacts breaks by mechanical loading, see criterion in Equation 23. Similarly, it has been demonstrated in the previous Section that grain reorganization occurs directly in the case $k_{0}^{init}\geq k_{0}^{attr}$, without the need for a notable reduction in cemented contacts. Indeed, the stable chain forces are predominant in this case. This is evidenced by the constant $k_{0}$ when the Young modulus $E_{m}$ remains constant during the debonding process, Figure 11b. As the Young modulus remains constant during the debonding, $k_{0}$ remains constant/granular reorganization does not occur. The cemented contacts break with a chemical rupture, as they are not loaded. The evolution of $k_{0}$ is a consequence of the softening of the grains. While the Young modulus decreases with the debonding process, the overlap between the grains increases, inducing grain reorganization (even if the particle positions remain largely unchanged) and a corresponding evolution of $k_{0}$.